OpenStudy (anonymous):

Jay has lost his mothers favorite necklace, so he will rent a metal detector to try to find it. A rental company charges a one-time rental fee of $15.00 plus $2.00 per hour to rent a metal detector. Jay has $35.00 to spend. What is the maximum amount of time he can rent the metal detector?

OpenStudy (anonymous):

One-time rental fee: $15 + $2/hour Total available money = $35 Rental time is x We have an inequality \[15 + 2x \leq 35\] \[\rightarrow x \leq (35 - 15) / 2 = 20 / 2 = 10\] So the maximum amount of time he can rent the metal detector is 10 hours

OpenStudy (anonymous):

ok then thanks